On July 1, 2016, Steelman Company acquired a new equipment for $157,000 and estimated it would have a useful life of 10 years and residual value of $7,500. At the beginning of 2019, the company decided that the equipment would be used for nine more years (including all of 2019), and at the end of this time its residual value would be only $825. On November 1, 2020, the equipment was sold for $82,000. The company uses the straight-line method of depreciation and closes its books on December 31.

Give the necessary journal entries for the acquisition, depreciation, and disposal of this asset for the years 2016, 2019, and 2020. (Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

July 1, 2016

Dec. 31, 2016

Dec. 31, 2019

Nov. 1, 2020

(To record depreciation expense)

Nov. 1, 2020

(To record sale of equipment)
Answer #1
Journal Entries
01-jul-16 Equipment a/c Dr 157000
To Cash/Bank a/c 157000
31-dec-16 Depreciation Expense a/c* Dr 7475
To Equipment a/c 7475
31-dec-19 Depreciation Expense a/c** Dr 13200
To Equipment a/c 13200
01-nov-20 Depreciation Expense a/c*** Dr 11000
To Equipment a/c 11000
01-Nov-20 Cash a/c 82000
Loss On Sale of equipment (Balancing figure) 13425
To Equipment a/c# 95425

*(157000-7500)/10=> 14950*6/12 = 7475

**Total Depreciation from 1/7/16 - 1/1/19 => 14950*2.5 years= 37375. So, wdv as on 1/1/19= 157000-37375 = 119625.

So, depreciation for remaining period = (119625-875)/9 = 13200.

***Depreciation upto the date of sale = 13200*10/12 = 11000.

# Value of equipment on 01-nov-20 = 157000 - 37375 - 13200 - 11000 = 95425